OpenStudy (anonymous):

Write a two-step equation that needs the steps of addition then division in order to be solved, and then solve it. Show your work.

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@Jennaluv

OpenStudy (anonymous):

@jonjenkins7653

OpenStudy (anonymous):

@MekaylaCox

OpenStudy (anonymous):

Sorry that came out wrong ill do it over

OpenStudy (anonymous):

3x - 6 = 15 3x - 6 + 6 =15 + 6 3x = 21 3x/3 = 21/3 x=7 Hope this helps :)

OpenStudy (anonymous):

thanks! @MekaylaCox

OpenStudy (anonymous):

Np